# An innovative clay metaBrick-based motif to enhance thermal and acoustic insulation

### v1

@@ Abstract
Metamaterials have gained popularity in recent years as a promising avenue for producing innovative materials with distinctive properties that offer unprecedented features. In this study, we address the thermal and acoustic challenges of building materials. We numerically and experimentally investigate a clay metaBrick, which is an innovative metamaterial design based on the existing hollow brick that has shown strong sound and thermal performances. We evaluate the acoustic and thermal characteristics of the clay metaBrick, including sound transmission and heat resistance, using the finite element method. Furthermore, we validate the results experimentally by investigating the behaviors of the wall built on metaBricks in two tests: sound and thermal. We contrast the findings of the metaBrick based wall with performances of a standard hollow brick wall. The metaBrick offers enhanced acoustic and thermal insulation properties compared to the standard brick, with a compressive strength above standards required in the building materials.
keywords:
Metamaterials; Helmholtz resonators; Clay metaBrick; Sound insulation; Thermal insulation
journal: Elsevier Journal
@@ 1 Introduction
Recently, in response to climate change and noise pollution, building materials with particular characteristics that may be employed to provide efficient heat and sound insulation have become necessary al2022built ; wang2022experimental ; anjum2022sustainable . These materials are crucial for decreasing energy consumption while enhancing comfort and safety in both residential and commercial environments gao2023comprehensive ; sutcu2014thermal ; fringuellino1999sound . Most studies have focused on improving the thermal and acoustic insulation of hollow clay bricks, which are the most commonly used construction materials due to their durability, low cost, and availability vijayan2021evaluation . In this regard, several approaches have been explored to increase the insulating properties of clay bricks. A common method is to add insulation materials, such as phase change materials, to the structure of the bricks to improve their thermal resistance principi2012thermal ; mahdaoui2021building ; aketouane2018energy ; chihab2022thermal ; bachir2023numerical ; hichem2013experimental ; luo2023numerical . Other research looked at the use of nanocrystalline aluminum sludge to increase the thermal efficiency of perforated clay bricks santos2015enhancement . Acoustically, several clay bricks with distinct hollow cores were analyzed jacqus2011homogenised . Furthermore, it has been reported that additional layers to the hollow bricks improve their acoustic performance souza2020impact . Another method is to change the structure of the bricks themselves, for example, by modifying the geometry of the air holes al2017effect . Besides these approaches, clay brick still has limitations in terms of both thermal and acoustic insulation properties. Therefore, additional advancement in the development of more efficient insulation techniques remains necessary.
@@
To control both phenomena, sound and heat, multi-functional metamaterials that combine these two physics are required. Metamaterials are man-made materials that have been developed to exhibit unique features that are not naturally disposable Kadic2019 ; gao2022acoustic . These materials have the potential to provide simultaneous heat and sound insulation, which can significantly improve energy efficiency and comfort in buildings and other applications. On the one hand, acoustic metamaterials are a type of metamaterials that are used to manipulate sound waves cummer2016controlling . In order to improve sound insulation, a variety of resonators with different characteristics are being studied, including phononic crystals with a Bragg-based bandgap liu2000locally ; al2017formation , free-standing mei2012dark ; yang2015subwavelength ; zhang2020light ; jang2022lightweight ; gu2022laminated , space coiled li2016acoustic ; cai2014ultrathin , Fabry-Perot wu2000profiled , quarter-wavelength shen2021acoustic , and Helmholtz resonatorsli2016sound ; jimenez2016ultra ; jimenez2017rainbow . By exploiting the special characteristics of acoustic metamaterials, it is feasible to design metamaterials that effectively block, absorb or deflect sound waves. On the other hand, thermal metamaterials have been designed to manipulate heat flow han2014full ; li2021transforming . These materials have unique thermal properties that enable unprecedented control over heat transfer. By engineering thermal metamaterials, it is possible to create materials that can efficiently trap or release heat, which makes them useful for applications such as thermal insulation zhang2023diffusion .
@@
In this paper, we suggest a new design for a hollow clay brick called a ”metaBrick,” where the word ”meta” refers to the concept of metamaterials. The metaBrick is built on Helmholtz resonators, which are designed to reduce the transmission of both sound and heat. Helmholtz resonators are more interesting for creating low-frequency sound insulation hoppen2023helmholtz , which is the straightforward reason why we use Helmholtz resonators. To evaluate the acoustic and thermal characteristics of the metaBrick, we combine the finite element approach with experiments. The results of the metaBrick are compared to those of the standard clay brick. The first section places a strong focus on the numerical performance analysis of the bricks. The second section is dedicated to empirically evaluating the effectiveness of a wall made of metaBricks, which involves measuring the walls’ thermal, acoustic and mechanical properties.
@@ 2 Characteristics of the metaBrick
To contribute to the development of the building materials, we have developed the hollow bricks into novel bricks called metaBricks, referring to the exotic behavior of metamaterials applied to the specific case of standard hollow bricks. The metaBrick design consists of an array of brick-shaped units with hollow cores, similar to a standard brick, but with the addition of slits at the top and the bottom of the standard hollow brick, as shown in Fig. 1. These features create Helmholtz resonators within the metaBrick, which can attenuate sound and reduce heat transfer. The dimensions of the metaBrick used in this study are $10$ $\mathrm{c}\mathrm{m}$ in width, $25$ $\mathrm{c}\mathrm{m}$ in length, and $20$ $\mathrm{c}\mathrm{m}$ in height. The slits on the top and bottom of the metaBrick are $0.5$ $\mathrm{c}\mathrm{m}$ in width and $25$ $\mathrm{c}\mathrm{m}$ in length. The metaBrick-based wall is illustrated in Fig. 1(a). The material parameters used in this study are depicted in Table 1.
Figure 1: Schematics of the Clay hollow metaBricks. (a) Wall made of clay metaBricks and mortar. (b) Clay hollow metabrick.
@@
Table 1: Building metaBrick mechanical and thermal properties
Material Young modulus Poisson’s Density Thermal conductivity Specific heat Sound speed
$\mathrm{G}\mathrm{P}\mathrm{a}$ ratio $\mathrm{k}\mathrm{g}\mathrm{/}\mathrm{m}^{3}$ $\mathrm{W}\mathrm{/}\mathrm{(}\mathrm{m}\mathrm{K}\mathrm{)}$ $\mathrm{J}\mathrm{/}\mathrm{(}\mathrm{k}\mathrm{g}\mathrm{K}\mathrm{)}$ $\mathrm{m}\mathrm{/}\mathrm{s}$
Clay 17 0.2 2000 0.5 900 -
Mortar 25 0.2 2300 1.8 880 -
Air - - 1.21 0.025 1005.3 343
@@ 2.1 Acoustic response
The numerical model used in the acoustic study is based on solving the acoustic wave propagation equations in both air and solid. Sound waves in air are governed by equation 1 for differential pressure $p$. The solid is considered as an isotropic linear elastic material, and the propagation of acoustic waves is determined by solving the weak formalism of the time-harmonic Navier equation (2).
.$\nabla.(\frac{1}{\rho_{air}}\nabla p)+\frac{\omega^{2}}{\rho_{air}C^{2}}p=0,$ (1)
@@
where $p$ is the acoustic pressure, u is the displacement vector, $\rho_{air}$ and $\rho_{s}$ are the densities of air and solid, respectively. $C$ is the speed of sound in air, $\omega$ is the angular frequency of the acoustic wave, $E$ is the Young’s modulus, and $\nu$ is the Poisson’s ratio.
In order to simulate the normal incident sound wave, a plane wave radiation boundary condition was utilized, and Perfectly Matched Layers (PMLs) were placed at both extremities of the medium to limit spurious reflections, as depicted in Fig. 2(a). Furthermore, the Floquet-Bloch periodic conditions are applied on the top and bottom of the bricks. We calculate the transmission in the frequency range ($20$ $\mathrm{H}\mathrm{z}$ to $5$ $\mathrm{k}\mathrm{H}\mathrm{z}$) for each brick, using equation (3).
$Transmission=20\times log_{10}(\frac{|p_{t}|}{|p_{i}|}),$ (3)
where $p_{i}$ and $p_{t}$ are the incident and transmitted pressures, respectively.
@@
Fig. 2(b) displays the sound transmission of two types of bricks, the standard brick (black curve) and the metaBrick (red curve), across a wide range of frequencies. It is evident that the standard brick has a higher sound transmission compared to the metaBrick, which is due to the tiny guide between the two successive bricks creating a Fabry-Perot resonator that has a resonance at $n\lambda/2$ (where $\lambda$ is the wavelength). The presence of slits, i.e., Helmholtz resonators, reduces the sound transmission by $20$ $\mathrm{d}\mathrm{B}$ across a broad frequency range from $500$ $\mathrm{H}\mathrm{z}$ to $3500$ $\mathrm{H}\mathrm{z}$, beyond which the limit of diffraction becomes apparent. The superior acoustic insulation properties of the metaBrick are clearly demonstrated through the transmission curve, as it effectively attenuates sound waves more than the classical brick. This finding, which applies without mortar, is of practical significance as it highlights the importance of reducing sound transmission through building materials to create acoustically comfortable structures that comply with regulatory standards.
@@
The response of the two types of bricks to plane wave excitation is further analyzed using screenshots shown in Fig. 3. The screenshots provide a visual representation of the bricks’ reactions to the plane wave at different frequencies and demonstrate that the metaBrick attenuates the sound wave more efficiently than the standard brick. This observation reinforces the conclusion that the metaBrick has superior acoustic insulation properties. The figure particularly highlights the energy trapping inside the Helmholtz resonators in the metabicks at different frequencies, unlike the classical ones, in which cavities play no particular role in the acoustic insulation. In summary, the study provides valuable insights into the acoustic performance of two types of brick and highlights the advantages of using the metaBrick for acoustic insulation applications.
@@ 2.2 Thermal response
For the thermal analysis, the heat transfer in solids and fluids equations was utilized to characterize thermal insulation through the bricks. By solving the transient heat equation (4), the temperature distribution within the unit cells is calculated. The methodology is based on applying a sinusoidal temperature within a $24$-hour period ranging between 00 and $50$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$ and determining the inner temperature of unit cells, as shown in Fig. 4(a). To accomplish this, the terminal boundaries of the configuration are exposed to ambient temperature, which is set to $T_{amb}=25$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$. The outer boundary of cells is vulnerable to the outer temperature $T_{out}(t)$, which describes the temperature variation over a $24$-hour period with the expression $T_{out}=25sin(\frac{2\pi}{24}t)+25$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$. Furthermore, the periodicity conditions are applied in the top and bottom of bricks.
@@
Fig. 4(b) reports a comparative analysis of the thermal properties of two types of bricks, namely the standard brick (curve in red) and the metaBrick (curve in blue). The results demonstrate that the metaBrick exhibits superior thermal insulation properties compared to the standard brick. Specifically, the metaBrick displays lower temperature values than the standard brick, indicating a greater capacity for retaining heat.
The superiority of the metaBrick’s thermal insulation properties is further supported by the screenshots provided in Fig. 5, which depict the temperature distribution across the bricks over time. The screenshots show that the metaBrick retains heat more effectively than the standard brick, thereby reducing the need for additional heating or cooling.
@@ 3 Characteristics of a metaBrick wall
Figure 6: Clay hollow metabrick experimental pictures, with a side view on the left and a top view on the right.
Furthermore, numerical simulations were carried out to study the behavior of both walls. The results obtained from the simulations were compared with the experimental data to validate the accuracy of the simulation models and gain further insight into the acoustic and thermal properties of the walls. Through the combined approach of experimental and numerical investigations, we were able to obtain a comprehensive understanding of the properties of the metaBrick wall and its advantages over standard brick walls. In Fig. 6, several images of the metaBrick used in our experimental tests are displayed. The images illustrate the top and side views of the metaBrick, side-by-side, allowing for a direct visual comparison of their geometrical attributes. The metaBrick’s unique design and composition can be observed, which include a mixture of various materials to enhance its thermal properties. These images provide a visual representation of the experimental setup and the materials used, which can help to understand the experimental results presented further down in this document.

### v2

# Clay metaBrick-based Motif to Enhance Thermal and Acoustic Insulation

@@ Abstract
Acoustic metamaterials have gained popularity as promising materials for enhancing noise reduction. Here, we explore the use of metamaterials, based on Helmholtz resonators (HRs), to enhance the performance of standard clay hollow brick. By incorporating HRs in the upper and lower hollows, we transform the standard brick into metaBrick, which is essentially designed based on metamaterial principles. We evaluate the acoustic and thermal performance of walls constructed with clay metaBricks, focusing on sound transmission loss and heat resistance. Both the finite element method and experimental analysis were employed to highlight the performance of metaBricks compared to standard clay bricks. Results show that metaBricks significantly enhance acoustic and thermal insulation, achieving an attenuation of $20$ $\mathrm{d}\mathrm{B}$ across a broad frequency range from $500$ to $2500$ $\mathrm{H}\mathrm{z}$ and an $8$% increase in thermal resistance. However, compressive strength is reduced by $33$%, though it remains above the standard requirements for building materials. These findings indicate that metaBrick is a promising building material, offering improved sound and thermal insulation.
Brahim Lemkalli1,2,∗, Saad Bensallam3, Muamer Kadic1, Sébastien Guenneau4,5, Hicham Jakjoud6, Abdellah Mir2, Younes Achaoui2
{affiliations}
1Université Marie et Louis Pasteur, Institut FEMTO-ST, CNRS, 25000 Besançon, France
2Laboratory of Optics, Information Processing, Mechanics, Energetics and Electronics, Department of Physics, Moulay Ismail University, B.P. 11201, Zitoune, Meknes, Morocco
3Hassania School of Public Works, Casablanca, Morocco
4The Blackett Laboratory, Physics Department, Imperial College London, SW7 2AZ, London, UK
5UMI 2004 Abraham de Moivre-CNRS, Imperial College London, SW7 2AZ, London, UK
6Laboratory of Energy Engineering, Materials and Systems, National Schools of Applied Sciences, Ibn Zohr University, Agadir, Morocco
lemkallibrahim@gmail.com
@@ 1 Introduction
Recently, in response to climate change and noise pollution, building materials with particular characteristics that may be employed to provide efficient heat and sound insulation have become necessary [1, 2, 3]. These materials are crucial for decreasing energy consumption while enhancing comfort and safety in both residential and commercial environments [4, 5, 6]. Most studies have focused on improving the thermal and acoustic insulation of hollow clay bricks, which are the most commonly used construction materials due to their durability, low cost, and availability [7]. In this regard, several approaches have been explored to increase the insulating properties of clay bricks. A common method is to add insulation materials, such as phase change materials, to the structure of the bricks to improve their thermal resistance [8, 9, 10, 11, 12, 13, 14]. Other research looked at the use of nanocrystalline aluminum sludge to increase the thermal efficiency of perforated clay bricks [15]. Acoustically, several clay bricks with distinct hollow cores were analyzed [16]. Furthermore, it has been reported that additional layers to the hollow bricks improve their acoustic performance [17]. Another method is to change the structure of the bricks themselves, for example, by modifying the geometry of the air holes [18]. Besides these approaches, clay brick still has limitations in terms of both thermal and acoustic insulation properties. Therefore, additional advancement in the development of more efficient insulation techniques remains necessary.
@@
To control both phenomena, sound and heat, multi-functional metamaterials that combine these two physics are required. Metamaterials are man-made materials that have been developed to exhibit unique features that are not naturally disposable [19, 20, 21, 22, 23, 24, 25]. These materials have the potential to provide simultaneous heat and sound insulation, which can significantly improve energy efficiency and comfort in buildings and other applications. On the one hand, acoustic metamaterials are a type of metamaterials that are used to manipulate sound waves [26, 27, 28, 29]. In order to improve sound insulation, a variety of resonators with different characteristics are being studied, including phononic crystals with a Bragg-based bandgap [30, 31], free-standing [32, 33, 34, 35, 36], space coiled [37, 38, 39], Fabry-Perot [40], quarter-wavelength [41], and Helmholtz resonators[42, 43, 44, 45, 46]. By exploiting the special characteristics of acoustic metamaterials, it is feasible to design metamaterials that effectively block, absorb or deflect sound waves. On the other hand, thermal metamaterials have been designed to manipulate heat flow [47, 48]. These materials have unique thermal properties that enable unprecedented control over heat transfer. By engineering thermal metamaterials, it is possible to create materials that can efficiently trap or release heat, which makes them useful for applications such as thermal insulation [49, 50].
Previous studies have also introduced the concepts of metawall and metabrick in the literature. For example, some works have focused on integrating porous materials into metabricks to enhance thermal performance [51]. Others have explored metawall unit bricks composed of rubber-metal-concrete composites [52]. However, the key challenge remains: how can we design a new type of metabrick that provides both thermal and acoustic insulation while relying on commonly used construction materials, as well as those already employed in industrial thermal and sound insulation systems?
@@
In this paper, we suggest a design for a hollow clay brick called a metaBrick, where the word ”meta” refers to the concept of metamaterials. The metaBrick is built on Helmholtz resonators, which are designed to reduce the transmission of both sound and heat. Helmholtz resonators are more interesting for creating low-frequency sound insulation [53], which is the straightforward reason why we use Helmholtz resonators. To evaluate the acoustic and thermal characteristics of the metaBrick, we combine the finite element method with experimental measurements. The results of the metaBrick are compared to those of the standard clay brick. The first section places a strong focus on the numerical performance analysis of the bricks. The second section is dedicated to empirically evaluating the effectiveness of a wall made of metaBricks, which involves measuring the walls’ thermal, acoustic and mechanical properties.
Figure 1: Schematics of the Clay hollow metaBricks. (a) Wall made of clay metaBricks and mortar. (b) Clay hollow metaBrick. (c) Clay hollow metaBrick experimental pictures, with a side view on the left and a top view on the right.
@@ 2 Characteristics of the metaBrick
To contribute to the development of the building materials, we have developed the hollow bricks into bricks based on metamaterials called metaBricks, referring to the exotic behavior of metamaterials applied to the specific case of standard hollow bricks. The metaBrick design consists of an array of brick-shaped units with hollow cores, similar to a standard brick, but with the addition of slits at the top and the bottom of the standard hollow brick, as shown in Figure 1. These features create Helmholtz resonators within the metaBrick, which can attenuate sound and reduce heat transfer. The dimensions of the metaBrick used in this study are $10$ $\mathrm{c}\mathrm{m}$ in width, $25$ $\mathrm{c}\mathrm{m}$ in length, and $20$ $\mathrm{c}\mathrm{m}$ in height. For the slits on the top and bottom of the metabrick, we conduct a series of simulations (see Figure 9) to study the effect of slit size on the sound transmission loss. Based on the results, we select slits with a width of $0.5$ $\mathrm{c}\mathrm{m}$ and a length of $25$ $\mathrm{c}\mathrm{m}$. The metaBrick-based wall is illustrated in Figure 1(a). The material parameters used in this study are depicted in Table 1.
@@
Table 1: Building metaBrick mechanical and thermal properties.
Material Clay Mortar Air
Young modulus ($\mathrm{G}\mathrm{P}\mathrm{a}$) 17 25 -
Poisson’s ratio 0.2 0.2 -
Density ($\mathrm{k}\mathrm{g}\mathrm{/}\mathrm{m}^{3}$) 2000 2300 1.21
Thermal conductivity ($\mathrm{W}\mathrm{/}\mathrm{(}\mathrm{m}\mathrm{K}\mathrm{)}$) 0.5 1.8 0.025
Specific heat ($\mathrm{J}\mathrm{/}\mathrm{(}\mathrm{k}\mathrm{g}\mathrm{K}\mathrm{)}$) 900 880 1005.3
Sound speed ($\mathrm{m}\mathrm{/}\mathrm{s}$) - - 343
@@ 2.1 Acoustic response
The numerical model used in the acoustic study is based on solving the acoustic wave propagation equations in both air and solid. Sound waves in air are governed by the Helmholtz equation 1 for differential pressure $p$, whereas acoustic waves in the isotropic linear elastic material are governed by the time-harmonic Navier equation (2):
$\nabla^{2}p+\frac{\omega^{2}}{C^{2}}p=0,$ (1)
@@
[…]
A multiphysics model is used to solve the acoustic wave propagation equations in both air and solids for the sound transmission analysis. As part of the multiphysics coupling, we use the following air-solid interface boundary conditions that couple the equations corresponding to waves propagating in both media[54]:
[…]
where $\sigma$ is the stress tensor for an isotropic linear elastic material, given by $\sigma=\frac{E}{2(1+\nu)}(\frac{1}{(1-2\nu)}\nabla.\textbf{u}+\nabla\textbf{u})$ and n is the outward unit normal to the air or solid.
@@
In order to simulate the normal incident sound wave, a plane wave radiation boundary condition was utilized, and Perfectly Matched Layers (PMLs) were placed at both extremities of the medium to limit spurious reflections, as depicted in Figure 2(a). Furthermore, the Floquet-Bloch periodic conditions are applied on the top and bottom of the bricks. For the meshing, we take into account a maximum element size of $\lambda_{\text{min}}/10$ and tetrahedral elements. We then calculate the transmission in the frequency range ($20$ $\mathrm{H}\mathrm{z}$ to $5$ $\mathrm{k}\mathrm{H}\mathrm{z}$) for each brick, using equation (5).
[…]
where $V_{1}$ and $V_{2}$ are the volumes of the media corresponding to the incident and transmitted waves, respectively, and $p_{i}$ and $p_{t}$ represent the incident and transmitted pressures.
Figure 2(b) displays the sound transmission of two types of bricks, the standard brick (black curve) and the metaBrick (red curve), across a wide range of frequencies. It is evident that the standard brick has a higher sound transmission compared to the metaBrick, which is due to the tiny guide between the two successive bricks creating a Fabry-Perot resonator that has a resonance at $n\lambda/2$ (where $\lambda$ is the wavelength). The presence of slits, i.e., Helmholtz resonators, reduces the sound transmission by $20$ $\mathrm{d}\mathrm{B}$ across a broad frequency range from $500$ $\mathrm{H}\mathrm{z}$ to $3500$ $\mathrm{H}\mathrm{z}$, beyond which the limit of diffraction becomes apparent. The superior acoustic insulation properties of the metaBrick are clearly demonstrated through the transmission curve, as it effectively attenuates sound waves more than the classical brick. This finding, which applies without mortar, is of practical significance as it highlights the importance of reducing sound transmission through building materials to create acoustically comfortable structures that comply with regulatory standards.
The response of the two types of bricks to plane wave excitation is further analyzed using snapshots of the acoustic pressure, as shown in Figure 3. For the corresponding displacement fields, see Figure 10. The screenshots provide a visual representation of the bricks’ reactions to the plane wave at different frequencies and demonstrate that the metaBrick attenuates the sound wave more efficiently than the standard brick. This observation reinforces the conclusion that the metaBrick has superior acoustic insulation properties. The figure particularly highlights the energy trapping inside the Helmholtz resonators in the metabicks at different frequencies, unlike the classical ones, in which cavities play no particular role in the acoustic insulation. In summary, the study provides valuable insights into the acoustic performance of two types of brick and highlights the advantages of using the metaBrick for acoustic insulation applications.
@@ 2.2 Thermal response
For the thermal analysis, the heat transfer in solids and fluids equations was utilized to characterize thermal insulation through the bricks. By solving the transient heat equation (6), the temperature distribution within the unit cells is calculated. The methodology is based on applying a sinusoidal temperature within a $24$-hour period ranging between 00 and $50$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$ and determining the inner temperature ($T_{inner}$) of unit cells, as shown in Figure 4(a). To accomplish this, the terminal boundaries of the configuration are exposed to ambient temperature, which is set to $T_{amb}=25$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$, which means that the inner temperature at the initial time equals to $T_{amb}$. The outer boundary of cells is vulnerable to the outer temperature $T_{outer}(t)$, which describes the temperature variation over a $24$-hour period with the expression $T_{outer}=25sin(\frac{2\pi}{24}t)+25$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$. Furthermore, the periodicity conditions are applied in the top and bottom of bricks.
@@
Figure 4(b) reports a comparative analysis of the thermal properties of two types of bricks, namely the standard brick (curve in red) and the metaBrick (curve in blue). The results demonstrate that the metaBrick exhibits superior thermal insulation properties compared to the standard brick. After $6$ hours of exposure to the sinusoidal outer temperature, the interior temperature of the metaBrick reaches a maximum of $27$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$, while the standard brick reaches $29$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$, resulting in a $2$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$ difference in the peak temperature and a $2$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$ difference in the lowest temperature. That means that the metabrick maintains a comfortable internal temperature, making it an excellent choice for thermal insulation over the standard one. The metaBrick, in particular, has lower temperature values than the standard clay brick, suggesting a stronger ability for retaining heat, which can act as an alternative solution to the greenhouse effect.
The superiority of the metaBrick’s thermal insulation properties is further supported by the screenshots provided in Figure 5, which depict the temperature distribution across the bricks over time. The screenshots show that the metaBrick retains heat more effectively than the standard brick, thereby reducing the need for additional heating or cooling.
@@ 3 Characteristics of a metaBrick wall
Furthermore, numerical simulations were carried out to study the behavior of both walls. The results obtained from the simulations were compared with the experimental data to validate the accuracy of the simulation models and gain further insight into the acoustic and thermal properties of the walls. Through the combined approach of experimental and numerical investigations, we were able to obtain a comprehensive understanding of the properties of the metaBrick wall and its advantages over standard brick walls. In Figure 1(c), several images of the metaBrick used in our experimental tests are displayed. The images illustrate the top and side views of the metaBrick, side-by-side, allowing for a direct visual comparison of their geometrical attributes. The metaBrick’s unique design and composition can be observed, which include a mixture of various materials to enhance its thermal properties. These images provide a visual representation of the experimental setup and the materials used, which can help to understand the experimental results presented further down in this document.